Andy and Ben took 4 hours to paint a room together. Andy could paint the same room on his own in 7 hours. How long would it take Ben to paint the same room on his own? ( Leave the answer as a mixed number )

Answer by stanbon(75887)

Andy and Ben took 4 hours to paint a room together. Andy could paint the same room on his own in 7 hours. How long would it take Ben to paint the same room on his own? ( Leave the answer as a mixed number )
---------------------------------------------
Together DATA:
time = 4 hrs/job ; rate = 1/4 job/hr
----------------
Andy DATA:
time = 7 hrs/job ; rate = 1/7 job/hr
-----------------
Ben DATA:
time = x hrs/job ; rate = 1/x job/hr
--------------------
EQUATION:
rate + rate = together rate
1/x+ 1/7 = 1/4
Multiply thru by 28x to get:
28 + 4x = 7x
3x = 28
x = 38/3 = 12 2/3 hrs (time for Ben to do the job alone)
================
Cheers,
Stan H.